Write an equation in slope intercept form:
The line passes through (2,4) and has the same y-intercept as line line whose equation is x - 4y = 8.

Answer by mananth(16949)

x - 4y = 8.
4y= x-8
y=x/4 -2
the y intercept is -2 (0,-2)
passing through (2,4)
x1 y1 x2 y2
0 -2 2 4

slope m = (y2-y1)/(x2-x1)
( 4 - -2 )/( 2 - 0 )
( 6 / 2 )
m= 3

Plug value of the slope and point ( 2 , 4 ) in
Y = m x + b
4.00 = 6 + b
b= 4 - 6
b= -2
So the equation will be
Y = 3 x -2
m.ananth@hotmail.ca